Parth (parthkohli):

Solve \(x(x + 3) = -9\) NOT by the quadratic formula.

OpenStudy (anonymous):

Well, clearly x cannot be positive, as that would mean x+3 is positive too. x must be negative, but no less than -3, or else x+3 would also be negative.

Parth (parthkohli):

Complex numbers?

Parth (parthkohli):

I know them by the quadratic formula

OpenStudy (anonymous):

maybe?

OpenStudy (anonymous):

Oh, right. I forgot to assume the real-ness of x... I'm lost :D

OpenStudy (phi):

you can complete the square as an alternative to the quadratric formula

OpenStudy (anonymous):

That's kinda like using the quadratic formula anyway? :D

OpenStudy (phi):

you will get complex roots

Parth (parthkohli):

Well\[x^2 + 3x + 2.75 = -9 + 2.75\]

OpenStudy (anonymous):

What's with the 2.75? You mean 2.25, right?

Parth (parthkohli):

Yes, whoops

OpenStudy (anonymous):

can i use compeleting square ??? or factor mehod ?

Parth (parthkohli):

\[(x + 1.5)^2 = -9 + 2.25\]Yes, I wanna know how to factor this over \(\mathbb C\) without knowing the roots!

OpenStudy (anonymous):

It seems like it's already almost done?

OpenStudy (anonymous):

well, i guess this is a good condidate for completing square .

Parth (parthkohli):

OK, completing the square worked out well. Thanks @phi :-)